OpenStudy (anonymous):

48,39,30,21, ___, ___ complete pattern

OpenStudy (kirbykirby):

12, 3

OpenStudy (kirbykirby):

You subtract 9 at every number 48 - 9 = 39 39 - 9 = 30 etc..

OpenStudy (kirbykirby):

30 - 9 = 21 21 - 9 = 12 12 - 9 = 3